A metal chloride dissolves endothermally and when 7.45g of its anhydrous form are dissolved in excess of water, the amount of heat absorbed is X kJ. Calculate the enthalpy of solution if molar mass of chlorine is 74.5

Grade:12

2 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Let metal chloride is MCl
If here molar mass of metal chloride is 74.5g/mol
so moles of metal chloride=7.45/74.5=0.1mol
as enthalpy of solution is given as heat absorbed/released per mole
so heat of solution(per mole)=x*10 kJ=10X kJ
shivi
7 years ago
Heat absorbed, q = −cm∆T where, c= specific heat capacity of water, 4.18 kJ kg−1 oC−1 m = mass of water in kg, ∆T=change in temperature in oC q = X kJ (given) Number of moles of the metal chloride = 7.4574.5mol= 0.1 molNow, the enthalpy of solution is heat absorbed per mole, i.e, ∆H = qn= X0.1= 10X kJ mole